# JUHAN ARU, PHILÉMON BORDEREAU

### v1

@@ 2.2. Constructing the non-crossing $\mathrm{SLE}_{4}$ measure
(13) $\frac{\mathrm{d}\tilde{\nu}_{D^{\prime}}}{\mathrm{d}\tilde{\nu}_{D}}(\eta)=\sum_{\begin{subarray}{c}\mathbf{b}=(\mathbf{b}_{k},+\lambda,\dots,+\lambda)\\ \mathbf{b}_{k}\in\left\{\pm\lambda\right\}^{k}\end{subarray}}\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}}}\frac{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}_{k}}}}{\mathrm{d}\mathbb{P}_{D}^{g_{\mathbf{b}}}}(\eta)\frac{Z_{D^{\prime},\mathbf{b}_{k}}}{Z_{D,\mathbf{b}}}=\sum_{\begin{subarray}{c}\mathbf{b}=(\mathbf{b}_{k},+\lambda,\dots,+\lambda)\\ \mathbf{b}_{k}\in\left\{\pm\lambda\right\}^{k}\end{subarray}}\mathbf{1}_{\eta\subset\mathcal{A}_{\mathbf{b}},\eta\subset D^{\prime}}\exp(\frac{1}{2}m_{D}(\eta,D\setminus D^{\prime})).$
@@
To carry out this computation, we use Lemma 10 with $\mathbb{P}_{X}=\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}}}$ (now abusing notation to drop the superscript $\mathbf{b}_{k}$), $\mathbb{P}_{Y_{1}}=\mathbb{P}_{D}^{g_{\mathbf{b}}}$, and $\mathbb{P}_{Y_{2}}$ corresponding to the law of $\xi:=T_{\mathcal{B}_{r}^{\prime}}\Phi_{D}^{0}$, the restriction of the GFF to $\mathcal{B}_{r}^{\prime}$ (recall that $\mathcal{B}_{r}^{\prime}=(\mathcal{B}^{\prime}\setminus\mathcal{B})\cap D$). As noted earlier, conditionally on $\eta\sim\mathbb{P}_{D}^{g_{\mathbf{b}}}$ and $\left\{\eta\subset D^{\prime}\right\}$, we have that $\xi\sim\mathcal{N}(0,G_{D\setminus\eta}|_{\mathcal{B}_{r}^{\prime}})$. We can then write
@@
Proposition 11.
For all $\xi\in H^{-1}(\mathcal{B}_{r}^{\prime})$, it holds that $\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}}}\ll\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}},\xi}$, and we have explicitly that
(14) $\frac{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}}}}{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}},\xi}}(\eta)=\mathbf{1}_{\eta\subset\mathcal{A}_{\mathbf{b}},\eta\in D^{\prime}}\exp(\frac{1}{2}(\xi,(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})\xi)_{\mathcal{B}_{r}^{\prime}}-2\lambda((H_{\partial D^{\prime}})_{\mathcal{B}_{\mathbf{b}}^{-}}^{\mathcal{B}_{r}^{\prime}}1_{\mathcal{B}_{\mathbf{b}}^{-}},\xi)_{\mathcal{B}_{r}^{\prime}}).$
We first pursue the proof of the theorem with this proposition in hand, and will turn to its proof just after. The computation of the expectation now relies on applying Lemma 6 to Equation (14).
To simplify the corresponding expression, first notice that, thanks to Lemma 1,
$I-(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})G_{D\setminus\eta}=I-H_{D^{\prime}}H_{D\setminus\eta}.$
We dropped the subscripts and superscripts indicating that the operators are seen on $H^{-1}(\mathcal{B}_{r}^{\prime})$, for more readability. Consequently, the determinant term in the expectation of Equation (14) simplifies to
$(\mathrm{det}_{F}(I-(G_{D\setminus\eta})^{1/2}(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})(G_{D\setminus\eta})^{1/2}))^{-1/2}$ $=(\mathrm{det}_{F}(I-(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})(G_{D\setminus\eta})))^{-1/2}$
@@
by Proposition 2 (note that this Proposition implicitly contains the statement that the determinant is well-defined, and hence that the use of Lemma 6 is legitimate).
On the other hand, the term corresponding to the exponential in Lemma 6 can be rewritten as
$\exp(\frac{1}{2}\norm{2\lambda(I-G_{D\setminus\eta}^{1/2}(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})G_{D^{\prime}\setminus\eta}^{1/2})^{-1/2}G_{D\setminus\eta}^{1/2}(H_{\partial D^{\prime}})_{\mathcal{B}_{\mathbf{b}}^{-}}^{\mathcal{B}_{r}^{\prime}}1_{\mathcal{B}_{\mathbf{b}}^{-}}}^{2})$
$=$ $\exp(2\lambda^{2}((H_{\partial D^{\prime}})_{\mathcal{B}_{\mathbf{b}}^{-}}^{\mathcal{B}_{r}^{\prime}}1_{\mathcal{B}_{\mathbf{b}}^{-}},G_{D\setminus\eta}(I-(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})G_{D\setminus\eta})^{-1}(H_{\partial D^{\prime}})_{\mathcal{B}_{\mathbf{b}}^{-}}^{\mathcal{B}_{r}^{\prime}}1_{\mathcal{B}_{\mathbf{b}}^{-}})_{\mathcal{B}_{r}^{\prime}})$
@@
(15) $=$ $\exp(2\lambda^{2}(1_{\mathcal{B}_{\mathbf{b}}^{-}},(H_{\partial D}-H_{\partial D^{\prime}})1_{\mathcal{B}_{\mathbf{b}}^{-}})_{\mathcal{B}_{\mathbf{b}}^{-}}).$
This can be re-written as a difference of Dirichlet energies, as in Section 1.3, so that we finally obtain
$\frac{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}}}}{\mathrm{d}\mathbb{P}_{D}^{g_{\mathbf{b}}}}=\mathbf{1}_{\eta\subset\mathcal{A}_{\mathbf{b}},\eta\subset D^{\prime}}\exp(\frac{1}{2}m_{D}(\eta,D\setminus D^{\prime}))\exp(\frac{1}{2}(\norm{g_{D,\mathbf{b}}}_{\nabla(D)}^{2}-\norm{g_{D^{\prime},\mathbf{b}}}_{\nabla(D^{\prime})}^{2})).$
Replacing this in Equation (13) with the right expression of $Z_{D,\mathbf{b}},Z_{D^{\prime},\mathbf{b}}$ finishes the proof. ∎
Let us mention as a remark that the reason we rely on induction is to avoid having to deal with both boundaries at once. It would be possible to consider $D^{\prime}$ a test domain, $\xi=(\xi_{l},\xi_{r})$ a Gaussian process on $\mathcal{B}_{l}^{\prime}\times\mathcal{B}_{r}^{\prime}$ with covariance matrix having diagonal values $(G_{D})_{\mathcal{B}_{l}^{\prime}}^{\mathcal{B}_{l}^{\prime}},(G_{D})_{\mathcal{B}_{r}^{\prime}}^{\mathcal{B}_{r}^{\prime}}$, and follow the same approach as above: however the expression (14) and the computation of the Gaussian integral that results become more involved.
We now turn to the proof of Proposition 11.
Proof of Proposition 11.
@@
where $u_{\xi}$ is the harmonic extension of $\xi$ to $D^{\prime}$. Note that $u_{\xi}^{\varepsilon}$ and $u_{\xi}$ coincide in $D^{\prime}_{\varepsilon}$. Let $\mathbb{Q}$ be the law of $\Phi:=\Phi_{D^{\prime}}^{0}+g_{\mathbf{b}}$ and ~$\tilde{\mathbb{Q}}_{\varepsilon}$ defined as
@@
so that $\tilde{\Phi}:=\Phi-u_{\xi}^{\varepsilon}$ is a zero-boundary GFF in $D^{\prime}$ under ~$\tilde{\mathbb{Q}}_{\varepsilon}$, by Girsanov. Let $\eta,\tilde{\eta}$ be the level-lines for $\Phi,\tilde{\Phi}+u_{\xi}$ respectively.
Like in [ALS22, Proposition 3.3], we argue that the events $\left\{\mathrm{dist}(\eta,\mathcal{B}_{r}^{\prime})>\varepsilon\right\},\left\{\mathrm{dist}(\tilde{\eta}_{\varepsilon},\mathcal{B}_{r}^{\prime})>\varepsilon\right\}$ are (~$\tilde{\mathbb{Q}}_{\varepsilon}$-)a.s. equal, and that on these events $\eta$ and $\tilde{\eta}_{\varepsilon}$ coincide. Whereas $(\eta,\Phi_{\eta})$ is a local set under $\mathbb{Q}$ by definition, $(\eta_{\tau},\Phi_{\eta_{\tau}})$ is also a local set [SS09, Lemma 3.12], where $\tau$ is the exit time of $\eta$ of $D_{\varepsilon}^{\prime}$, and $\Phi_{\eta_{\tau}}$ is the harmonic function in $D^{\prime}\setminus\eta_{\tau}$ with boundary values given by $-\lambda$ (resp. $+\lambda$) on the left (resp. right) boundary of $\eta_{\tau}$, and coinciding with $g_{\mathbf{b}}$ on the other parts of $\partial D^{\prime}$. Now by Lemma 5, $(\eta_{\tau},\tilde{\Phi}_{\eta_{\tau}})$ is also a local set under ~$\tilde{\mathbb{Q}}_{\varepsilon}$, with $\tilde{\Phi}_{\eta_{\tau}}=\Phi_{\eta_{\tau}}-(u_{\xi}^{\varepsilon})_{\eta_{\tau}}$. On the other hand, $(\tilde{\eta}_{\tau},\tilde{\Phi}_{\eta_{\tau}})$, the level-line for $\tilde{\Phi}+u_{\xi}$, is also a local set under ~$\tilde{\mathbb{Q}}_{\varepsilon}$. Furthermore, the fields $\Phi$ and $\tilde{\Phi}+u_{\xi}$ coincide on $D_{\varepsilon}^{\prime}$. It follows from local uniqueness of level-lines ([ASW19, Lemma 15]) that $\eta_{\tau}=\tilde{\eta}_{\tau}$$\tilde{\mathbb{Q}}_{\varepsilon}$-almost surely. If $\tau=+\infty$, i.e. $\eta\subset D_{\varepsilon}^{\prime}$, then it is clear that $\tilde{\eta}\subset D_{\varepsilon}^{\prime}$ and $\eta=\tilde{\eta}$.
This implies that the law $\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}},\xi}$ of the level line of $\Phi+u_{\xi}$ under ~$\tilde{\mathbb{Q}}_{\varepsilon}$, on the event that it stays at a distance $>\varepsilon$ from $\mathcal{B}_{r}^{\prime}$, is absolutely continuous w.r.t. the law $\mathbb{P}_{D^{\prime}}$ of the level line of $\Phi$ under $\mathbb{Q}$ on the event that it too stays at a distance $>\varepsilon$ from $\mathcal{B}_{r}^{\prime}$, and the Radon-Nikodym derivative is given the conditional expectation of the RHS in Equation (16) given $\eta$, which by Lemma 5 is given by:
@@
We stress here that the Lemma above applies to a zero-boundary GFF, and a local set with $\Phi_{\eta}=h_{\eta}=0$ on the boundary. In our framework, that amounts to considering the curve $\eta$ coupled with $\Phi-g_{\mathbf{b}}=\Phi_{D^{\prime}}^{0}$, and where the harmonic part $h_{\eta}$ is the difference of the two-valued function that is equal to $-\lambda$ on the left of $\eta$ and $+\lambda$ on the right, and of the harmonic function $g_{\mathbf{b}}$. We compute this expression on the event that $\mathrm{dist}(\eta,\mathcal{B}_{r}^{\prime})>\varepsilon$, and take $\varepsilon$ to 00. Since $u_{\xi}$ is harmonic in $D^{\prime}$, we can write
@@
using that $h_{\eta}$ is harmonic and zero on $\partial D^{\prime}$, and $u_{\xi}^{\varepsilon}-u_{\xi}$ vanishes on $\partial(D^{\prime}\setminus D^{\prime}_{\varepsilon})$. Note again the abuse of notation: the above denotes the distributional pairing between $\xi\in H^{-1}(\mathcal{B}_{r}^{\prime})$ and $\partial_{\nu}h_{\eta}\in H^{1}(\mathcal{B}_{r}^{\prime})$. The explicit expression for $h_{\eta}$ (or for $g_{\mathbf{b}}$) in terms of the Poisson kernel allows to further write
@@
(19) $((u_{\xi}^{\varepsilon})_{\eta},(u_{\xi}^{\varepsilon})_{\eta})_{\nabla(D^{\prime})}=(\xi,(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})\xi)_{\mathcal{B}_{r}^{\prime}}$
Notice that this expression makes sense as $H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)}$ is a smoothing operator, and that it suffices to prove it for $\xi$ smooth as for a general $\xi\in H^{-1}(\mathcal{B}_{r}^{\prime})$, taking $(\xi_{n})_{n\geq 0}\subset C_{0}^{\infty}(\mathcal{B}_{r}^{\prime})$ converging to $\xi$ in $H^{-1}(\mathcal{B}_{r}^{\prime})$, we can write Equation (19) for $\xi_{n}$ and take the limit $n\to\infty$. This follows from $H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)}$ being continuous on $H^{-1}(D^{\prime})$ and by $(u_{\xi_{n}}^{\varepsilon})_{\eta}$ converging to $(u_{\xi}^{\varepsilon})_{\eta}$ in $H_{0}^{1}(D^{\prime})$. Let $v_{\eta}$ be defined as
@@
(21) $(v_{\eta},v_{\eta})_{\nabla(D^{\prime})}-(u_{\xi},u_{\xi})_{\nabla(D^{\prime})}=\int_{\mathcal{B}_{r}^{\prime}}(v_{\eta}\cdot\partial_{\nu}v_{\eta}-u_{\xi}\cdot\partial_{\nu}u_{\xi}).$

### v2

@@ 2.2.2. Proof of the theorem
(17) $\frac{\mathrm{d}\tilde{\nu}_{D^{\prime}}}{\mathrm{d}\tilde{\nu}_{D}}(\eta)=\sum_{\begin{subarray}{c}\mathbf{b}=(\mathbf{b}_{k},+\lambda,\dots,+\lambda)\\ \mathbf{b}_{k}\in\left\{\pm\lambda\right\}^{k}\end{subarray}}\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}}}\frac{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}_{k}}}}{\mathrm{d}\mathbb{P}_{D}^{g_{\mathbf{b}}}}(\eta)\frac{Z_{D^{\prime},\mathbf{b}_{k}}}{Z_{D,\mathbf{b}}}=\sum_{\begin{subarray}{c}\mathbf{b}=(\mathbf{b}_{k},+\lambda,\dots,+\lambda)\\ \mathbf{b}_{k}\in\left\{\pm\lambda\right\}^{k}\end{subarray}}\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}},\eta\subset D^{\prime}}\exp(\frac{1}{2}m_{D}(\eta,D\setminus D^{\prime})).$
@@
To carry out this computation, we use Lemma 2.8 with $\mathbb{P}_{X}=\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}}}$ (now abusing notation to drop the superscript $\mathbf{b}_{k}$), $\mathbb{P}_{Y_{1}}=\mathbb{P}_{D}^{g_{\mathbf{b}}}$, and $\mathbb{P}_{Y_{2}}$ corresponding to the law of $\xi:=T_{\mathcal{B}_{r}^{\prime}}\Phi_{D}^{0}$, the restriction of the GFF to $\mathcal{B}_{r}^{\prime}$. As noted earlier, conditionally on $\eta\sim\mathbb{P}_{D}^{g_{\mathbf{b}}}$ and $\left\{\eta\subset D^{\prime}\right\}$, we have that $\xi\sim\mathcal{N}(0,G_{D\setminus\eta}|_{\mathcal{B}_{r}^{\prime}})$. We can then write
@@
Proposition 2.9.
For all fixed $\xi\in H^{-1}(\mathcal{B}_{r}^{\prime})$, it holds that $\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}}}\ll\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}},\xi}$, and we have explicitly that
(18) $\frac{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}}}}{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}},\xi}}(\eta)=\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}},\eta\subset D^{\prime}}\exp(\frac{1}{2}(\xi,(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})_{\mathcal{B}_{r}^{\prime}}^{\mathcal{B}_{r}^{\prime}}\xi)_{\mathcal{B}_{r}^{\prime}}-2\lambda((H_{\partial D^{\prime}})_{\mathcal{B}_{\mathbf{b}}^{-}}^{\mathcal{B}_{r}^{\prime}}1_{\mathcal{B}_{\mathbf{b}}^{-}},\xi)_{\mathcal{B}_{r}^{\prime}}).$
We first pursue the proof of Theorem 2.6 with this proposition in hand, and will turn to its proof just after. The computation of the expectation now relies on applying Lemma 2.3 to Equation (18). The determinant term reads [^5] In this proof all occurrences of the Green’s function $G_{D\setminus\eta}$ should be understood as the operator $(G_{D\setminus\eta})_{\mathcal{B}_{r}^{\prime}}^{\mathcal{B}_{r}^{\prime}}$ on $L^{2}(\mathcal{B}_{r}^{\prime})$, similarly for $G_{D^{\prime}\setminus\eta}$ and $G_{D}$: we drop the subscripts and superscripts to lighten the notation.
$(\mathrm{det}_{F}(I_{\mathcal{B}_{r}^{\prime}}$ $-(G_{D\setminus\eta})^{1/2}(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})_{\mathcal{B}_{r}^{\prime}}^{\mathcal{B}_{r}^{\prime}}(G_{D\setminus\eta})^{1/2}))^{-1/2}$
$=(\mathrm{det}_{F}(I_{\mathcal{B}_{r}^{\prime}}-(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})_{\mathcal{B}_{r}^{\prime}}^{\mathcal{B}_{r}^{\prime}}G_{D\setminus\eta}))^{-1/2}$
@@
by Proposition 1.2 (note that this proposition implicitly contains the statement that the determinant is well-defined, and hence that the use of Lemma 2.3 is legitimate).
On the other hand, the term corresponding to the exponential in Lemma 2.3 can be rewritten as
$\exp(\frac{1}{2}\norm{2\lambda(I_{\mathcal{B}_{r}^{\prime}}-G_{D\setminus\eta}^{1/2}(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})_{\mathcal{B}_{r}^{\prime}}^{\mathcal{B}_{r}^{\prime}}G_{D^{\prime}\setminus\eta}^{1/2})^{-1/2}G_{D\setminus\eta}^{1/2}(H_{\partial D^{\prime}})_{\mathcal{B}_{\mathbf{b}}^{-}}^{\mathcal{B}_{r}^{\prime}}1_{\mathcal{B}_{\mathbf{b}}^{-}}}^{2})$
$=$ $\exp(2\lambda^{2}((H_{\partial D^{\prime}})_{\mathcal{B}_{\mathbf{b}}^{-}}^{\mathcal{B}_{r}^{\prime}}1_{\mathcal{B}_{\mathbf{b}}^{-}},(I_{\mathcal{B}_{r}^{\prime}}-(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})_{\mathcal{B}_{r}^{\prime}}^{\mathcal{B}_{r}^{\prime}}G_{D\setminus\eta})^{-1}G_{D\setminus\eta}(H_{\partial D^{\prime}})_{\mathcal{B}_{\mathbf{b}}^{-}}^{\mathcal{B}_{r}^{\prime}}1_{\mathcal{B}_{\mathbf{b}}^{-}})_{\mathcal{B}_{r}^{\prime}})$
@@
(19) $=$ $\exp(2\lambda^{2}(1_{\mathcal{B}_{\mathbf{b}}^{-}},(H_{\partial D}-H_{\partial D^{\prime}})_{\mathcal{B}_{\mathbf{b}}^{-}}^{\mathcal{B}_{\mathbf{b}}^{-}}1_{\mathcal{B}_{\mathbf{b}}^{-}})_{\mathcal{B}_{\mathbf{b}}^{-}}),$
where we used the identities in Lemma 1.1. This can finally be re-written as a difference of Dirichlet energies, as in Section 1.3, so that we finally obtain
$\frac{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}}}}{\mathrm{d}\mathbb{P}_{D}^{g_{\mathbf{b}}}}=\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}},\eta\subset D^{\prime}}\exp(\frac{1}{2}m_{D}(\eta,D\setminus D^{\prime}))\exp(-\frac{1}{2}(\norm{g_{D,\mathbf{b}}}_{\nabla(D),\mathrm{reg}}^{2}-\norm{g_{D^{\prime},\mathbf{b}}}_{\nabla(D^{\prime}),\mathrm{reg}}^{2})).$
The exponential term in the equation above corresponds to $Z_{D,\mathbf{b}}/Z_{D^{\prime\prime}}$, which concludes the proof. ∎
We now turn to the proof of Proposition 2.9.
Proof of Proposition 2.9.
@@
where $u_{\xi}$ is the harmonic extension of $\xi$ to $D^{\prime}$. Note that $u_{\xi}^{\varepsilon}$ and $u_{\xi}$ coincide in $D^{\prime}_{\varepsilon}$. Let $\mathbb{Q}$ be the law of $\Phi:=\Phi_{D^{\prime}}^{0}+g_{\mathbf{b}}$ and $\tilde{\mathbb{Q}}_{\varepsilon}$ defined as
@@
so that $\tilde{\Phi}:=\Phi-u_{\xi}^{\varepsilon}$ is a zero-boundary GFF in $D^{\prime}$ under $\tilde{\mathbb{Q}}_{\varepsilon}$, by Girsanov. Let $\eta,\tilde{\eta}$ be the level-lines for $\Phi,\tilde{\Phi}+u_{\xi}$ respectively.
Like in [3, Proposition 3.3], we argue that the events $\left\{\mathrm{dist}(\eta,\mathcal{B}_{r}^{\prime})>\varepsilon\right\},\left\{\mathrm{dist}(\tilde{\eta},\mathcal{B}_{r}^{\prime})>\varepsilon\right\}$ are ($\tilde{\mathbb{Q}}_{\varepsilon}$-)a.s. equal, and that on these events $\eta$ and $\tilde{\eta}$ coincide. Whereas $(\eta,\Phi_{\eta})$ is a local set under $\mathbb{Q}$ by definition, $(\eta_{\tau},\Phi_{\eta_{\tau}})$ is also a local set [34, Lemma 3.12], where $\tau$ is the exit time of $\eta$ of $D_{\varepsilon}^{\prime}$, and $\Phi_{\eta_{\tau}}$ is the harmonic function in $D^{\prime}\setminus\eta_{\tau}$ with boundary values given by $\lambda$ (resp. $\lambda$) on the left (resp. right) boundary of $\eta_{\tau}$, and coinciding with $g_{\mathbf{b}}$ on the other parts of $\partial D^{\prime}$. Now by Lemma 2.2, $(\eta_{\tau},\tilde{\Phi}_{\eta_{\tau}})$ is also a local set under $\tilde{\mathbb{Q}}_{\varepsilon}$, with $\tilde{\Phi}_{\eta_{\tau}}=\Phi_{\eta_{\tau}}-(u_{\xi}^{\varepsilon})_{\eta_{\tau}}$. On the other hand, $(\tilde{\eta}_{\tau},\tilde{\Phi}_{\eta_{\tau}})$, the level-line for $\tilde{\Phi}+u_{\xi}$, is also a local set under $\tilde{\mathbb{Q}}_{\varepsilon}$. Furthermore, the fields $\Phi$ and $\tilde{\Phi}+u_{\xi}$ coincide on $D_{\varepsilon}^{\prime}$. It follows from local uniqueness of level-lines ([4, Lemma 15]) that $\eta_{\tau}=\tilde{\eta}_{\tau}$ $\tilde{\mathbb{Q}}_{\varepsilon}$-almost surely. If $\tau=+\infty$, i.e. $\eta\subset D_{\varepsilon}^{\prime}$, then it is clear that $\tilde{\eta}\subset D_{\varepsilon}^{\prime}$ and $\eta=\tilde{\eta}$.
This implies that the law $\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}},\xi}$ of the level line of $\Phi+u_{\xi}$ under $\tilde{\mathbb{Q}}_{\varepsilon}$, on the event that it stays at a distance $\varepsilon$ from $\mathcal{B}_{r}^{\prime}$, is absolutely continuous w.r.t. the law $\mathbb{P}_{D^{\prime}}$ of the level line of $\Phi$ under $\mathbb{Q}$ on the event that it too stays at a distance $\varepsilon$ from $\mathcal{B}_{r}^{\prime}$, and the Radon-Nikodym derivative is given the conditional expectation of the RHS in Equation (20) given $\eta$, which by Lemma 2.2 is given by:
@@
We stress here that Lemma 2.2 above applies to a zero-boundary GFF, and a local set with $\Phi_{\eta}=h_{\eta}=0$ on the boundary. In our framework, that amounts to considering the curve $\eta$ coupled with $\Phi-g_{\mathbf{b}}=\Phi_{D^{\prime}}^{0}$, and where the harmonic part $h_{\eta}$ is the difference of the two-valued function that is equal to $\lambda$ on the left of $\eta$ and $\lambda$ on the right, and of the harmonic function $g_{\mathbf{b}}$. We compute this expression on the event that $\mathrm{dist}(\eta,\mathcal{B}_{r}^{\prime})>\varepsilon$, and take $\varepsilon$ to 0. Since $u_{\xi}$ is harmonic in $D^{\prime}$, we can write
@@
using that $h_{\eta}$ is harmonic and zero on $\partial D^{\prime}$, and $u_{\xi}^{\varepsilon}-u_{\xi}$ vanishes on $\partial(D^{\prime}\setminus D^{\prime}_{\varepsilon})$. Note again the abuse of notation: the above denotes the distributional pairing between $\xi$ and $\partial_{\nu}h_{\eta}$. The explicit expression for $h_{\eta}$ (or for $g_{\mathbf{b}}$) in terms of the Poisson kernel allows to further write
@@
(23) $((u_{\xi}^{\varepsilon})_{\eta},(u_{\xi}^{\varepsilon})_{\eta})_{\nabla(D^{\prime})}=(\xi,(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})_{\mathcal{B}_{r}^{\prime}}^{\mathcal{B}_{r}^{\prime}}\xi)_{\mathcal{B}_{r}^{\prime}}$
Notice that this expression makes sense as $(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})_{\mathcal{B}_{r}^{\prime}}^{\mathcal{B}_{r}^{\prime}}$ is a smoothing operator, and that it suffices to prove it for $\xi$ smooth as for a general $\xi\in H^{-1}(\mathcal{B}_{r}^{\prime})$, taking $(\xi_{n})_{n\geq 0}\subset C_{0}^{\infty}(\mathcal{B}_{r}^{\prime})$ converging to $\xi$ in $H^{-1}(\mathcal{B}_{r}^{\prime})$, we can write Equation (23) for $\xi_{n}$ and take the limit $n\to\infty$. This follows from $H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)}$ being continuous on $H^{-1}(\mathcal{B}_{r}^{\prime})$ and by $(u_{\xi_{n}}^{\varepsilon})_{\eta}$ converging to $(u_{\xi}^{\varepsilon})_{\eta}$ in $H_{0}^{1}(D^{\prime})$. Let $v_{\eta}$ be defined as
@@
(25) $(v_{\eta},v_{\eta})_{\nabla(D^{\prime})}-(u_{\xi},u_{\xi})_{\nabla(D^{\prime})}=\int_{\mathcal{B}_{r}^{\prime}}(u_{\xi}\cdot\partial_{\nu}u_{\xi}-v_{\eta}\cdot\partial_{\nu}v_{\eta}).$
@@
2.2.3. Further remarks
In general, we know no closed form expression for the probability terms $\mathbb{P}_{D}^{g_{\mathbf{b}}}(\mathcal{A}_{\mathbf{b}})$. However, in the case of the annulus such an expression can be found via stochastic calculus. Let us explain its origins in the following remark.